OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $0.88, and a college-ruled notebook for $2.38. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $59.00. How many of each type were sold?

OpenStudy (anonymous):

What are your givens?

OpenStudy (anonymous):

List them out first.

OpenStudy (anonymous):

graph paper=40, college ruled=10

OpenStudy (anonymous):

g=graph paper, c=college-ruled. $0.88g+$2.38c=$59.00 (a total of $59.00 was made) g+c=50 (a total of 50 was sold) use substitution method. c=50-g 0.88g+2.38(50-5)=59 0.88g+119-2.38g=59 -1.5g+119=59 -1.5g=-60 (-1.5/-1,5)g=-60/-15 g=40 40+c=50 -40 -40 c=10

OpenStudy (anonymous):

Substitution is unnecessary and a complete waste of time in this situation. Just eliminate.

OpenStudy (anonymous):

can you show me? petewe

OpenStudy (anonymous):

and thank you nahomie

OpenStudy (anonymous):

Let x = graph let y = college 0.88x + 2.38y = 59 x + y = 50 What do you from there if you want to eliminate a variable?

OpenStudy (anonymous):

add/subtract that variable so they canel

OpenStudy (anonymous):

Yes but because the variables in each equation have different coefficients, how do we make it so that at least 1 variable in each equation is equal?

OpenStudy (anonymous):

that is what has me confused..i dont know

OpenStudy (anonymous):

If you change one side of an equation, what must be done to the other?

OpenStudy (anonymous):

you have to change that side too

OpenStudy (anonymous):

Exactly. Given that, how can we make the x in the second equation equal to the x in the first equation? 0.88x + 2.38y = 59 x + y = 50

OpenStudy (anonymous):

add them

OpenStudy (anonymous):

No. What do you have to do x to make it 0.88x?

OpenStudy (anonymous):

multiply???

OpenStudy (anonymous):

Good, now if you multiply x, what do you have to do to everything else in the equation?

OpenStudy (anonymous):

multiply as well

OpenStudy (anonymous):

Exactly, show me what you get.

OpenStudy (anonymous):

0.88x+2.09=51.92

OpenStudy (anonymous):

Good, now the x in both equations are equal. From here, subtract the equations and the x will cancel out. Isolate and solve for y. Once you have y, sub that into x + y = 50 to find x. Hope I helped.

OpenStudy (anonymous):

you did THANK YOU!